# AuthPrint: Fingerprinting Generative Models Against Malicious Model Providers

### v1

@@ Appendix D Fingerprint Recovery Attacks
$\Delta\mathrm{MSE}(x^{\prime},x)=\frac{1}{l}\sum_{i\in S}\left[(r_{i}-x^{\prime}_{i})^{2}-(r_{i}-x_{i})^{2}\right].$
Only manipulated fingerprint pixels contribute to the delta in MSE, therefore we obtain:
$\Delta\mathrm{MSE}(x^{\prime},x)=\frac{1}{l}\sum_{i\in M\cap S}\left[(r_{i}-x^{\prime}_{i})^{2}-(r_{i}-x_{i})^{2}\right].$
Assuming that $R$ is accurate, i.e., $r_{i}\approx x_{i}$, we obtain:
$\Delta\mathrm{MSE}(x^{\prime},x)=\frac{1}{l}\sum_{i\in M\cap S}\left[(x_{i}-x^{\prime}_{i})^{2}\right],$
thus, we can approximate the delta in MSE as:
$\Delta\mathrm{MSE}(x^{\prime},x)\approx\frac{|M\cap S|\cdot\delta^{2}}{l}.$
Thus, the detector rejects if
$|M\cap S|\geq k:=\left\lceil\frac{\tau l}{\delta^{2}}\right\rceil,$
where $k$ is the minimum number of manipulated fingerprint pixels required to trigger detection.
As an example, suppose $l=32$, $\tau=0.01$, and $\delta=0.5$. Then $\delta^{2}=0.25$, and the rejection threshold becomes:
$k=\left\lceil\frac{0.01\times 32}{0.25}\right\rceil=\left\lceil 1.28\right\rceil=2.$
This means the detector will reject any image where at least $k=2$ fingerprint pixels in $S$ have been manipulated.
@@
Theoretical Query Complexity Estimation by NATGT.
We reduce Step 1 of the problem of fingerprint recovery to a group testing task. Under the non-adaptive threshold group testing (NATGT) framework, the query complexity required to recover $S$ depends on the fingerprint length $l$, the rejection threshold $k$, and the image dimensionality $d$.
Theoretical analysis of NATGT shows that the query complexity in our problem satisfies:
@@
as derived in (Bui, Cheraghchi, and Echizen 2021). Their analysis considers NATGT in the general case with a gap, where the test outcome may be uncertain if the number of intersecting items lies between two thresholds. In our fingerprint recovery setting, however, we consider the gap-free regime, where tests yield strictly binary outcomes (accept or reject) with only one threshold $k$. By setting the gap to zero in their results and substitute with our parameters in the problem, we obtain the above complexity estimation in Equation 1.
Impact of $k/l$ on Query Complexity.
To understand the relationship between AuthPrint parameters and query complexity estimation under NATGT, we examine how the query complexity $T$ varies with respect to the length ratio $k/l$ according to Equation 1. We systematically evaluate this relationship by fixing different fingerprint lengths $l\in\{32,64,128,256,512\}$ and computing the corresponding query complexity as $k/l$ ranges from $1/32$ to $1/2$. We maintain $d=3\times 256^{2}$ for all data points. The results are shown in Figure 10.
Figure 10: Query complexity $T$ plotted against length ratio $k/l$ under different fixed fingerprint lengths $l$. All experiments use image dimensionality $d=3\times 256^{2}$.
Reconstructor Stability Experiment.
We validate the assumption that the reconstructor output $R(x)$ remains stable under small manipulations to image pixels for images from the authentic model used for training $R$. Using 100 generated images from a pre-trained StyleGAN2 model (FFHQ, 70k train data, ADA augmentation), we apply 1,000 random manipulation trials per image, varying the number of manipulated pixels $m\in\{1,2,4,\ldots,65536\}$. The reconstructor $R$ is trained with fingerprint length $l=32$. In each trial, we measure the mean $\ell_{1}$ distance between the reconstructor outputs on the clean and manipulated images. The results are shown in Figure 11.
Figure 11: Reconstructor stability under pixel manipulation. Mean $\ell_{1}$ distance between reconstructor outputs remains low for small $m$, supporting the stability assumption.
Appendix E Related Work
Cryptographic Approaches. Ensuring that an output originates from a specific ML model is an emerging challenge. Most cryptographic approaches focus on proofs for the correctness of the model’s computation. Early probabilistic proof systems for verifiable computation were highly interactive (Wahby et al. 2017; Ghodsi, Gu, and Garg 2017), which renders them unsuitable to our setting. Functional commitments offer zero-knowledge (ZK) proofs of the evaluation of a committed function for functions that can be represented as arithmetic circuits  (Boneh, Nguyen, and Ozdemir 2021). Subsequent works in the ZK-ML literature have implemented small neural networks as arithmetic circuits, enabling zk-SNARKs (Kang et al. 2023; Lee et al. 2024). However, the conversion to an arithmetic circuit has a cost in model accuracy that would discourage adoption.
A different approach is to create an immutable snapshot of the model by loading it into a trusted execution environment (TEE) (Tramer and Boneh 2018). The auditor can then publish a certificate with the TEE’s public key that clients may use to verify through remote attestation. TEEs on GPUs are still under development, and thus most model providers following this approach would not be able to benefit from the speedups of GPU computing for popular large generative models. To address this issue, Tramèr and Boneh propose delegating computations to an untrusted GPU (Tramer and Boneh 2018) and verify them with a probabilistic integrity check. To provide confidentiality toward the GPU, they use a symmetric cipher, which requires quantizing the model weights, thus this approach not only impacts latency but also model accuracy.
The practical limitations of cryptographic methods highlight the need for lightweight approaches that trade verification guarantees for practical feasibility.
Other Related Work on Fingerprinting Techniques. Fingerprinting techniques like the ones used for model fingerprinting have also been explored for deepfake vs. real detection. For example, Nataraj et al. observed that GANs tend to exhibit higher co-occurrence of pixels across generated images than real images do, and train a classifier to distinguish between GAN and real based on co-occurrence matrices (Nataraj et al. 2019). Although these techniques share similarities, they address a binary classification problem, grouping all models into the same “artificial” class.
In addition, the term model fingerprinting has been used to describe techniques that embed an identifier in the model’s parameters with the goal of detecting unauthorized uses of the model (Chen, Rohani, and Koushanfar 2018). Although these methods share the objective of protecting intellectual property with the model fingerprinting methods we build upon, the underlying techniques are differ fundamentally and operate under different threat models: the verifier requires white-box access to the model in order to verify information encoded in the weights.

### v2

@@ V-F2 Fingerprint Recovery Attacks
$\Delta\mathrm{MSE}(x^{\prime},x)=\frac{1}{l}\sum_{i\in M\cap S}\left[(x_{i}-x^{\prime}_{i})^{2}\right]\approx\frac{|M\cap S|\cdot\delta^{2}}{l}.$ (2)
The detector rejects only if the change in MSE exceeds a fixed threshold $\tau$. Thus, the minimum number of manipulated fingerprint pixels needed for rejection is:
$|M\cap S|\geq k:=\left\lceil\frac{\tau l}{\delta^{2}}\right\rceil.$ (3)
This quantity $k$ defines the rejection threshold—the smallest overlap size between $M$ and $S$ needed to trigger rejection. As a representative example from our evaluation, with $l=32$, $\tau=0.01$, and $\delta^{2}=0.25$, the calculation yields $k=\lceil 1.28\rceil=2$.
The problem of learning a subset of targeted pixels by systematically querying subsets of pixels and receiving one bit of information if the subset contains at least $k$ targeted pixels mirrors the non-adaptive threshold group testing (NATGT) problem [30]: the secret fingerprint set $S$ corresponds to the hidden “defective” set in NATGT, and each attacker query corresponds to NATGT’s test set $M_{i}$. The detector returns a binary signal—accept or reject—based on whether $|M_{i}\cap S|\geq k$, where $k$ is known as the threshold. In NATGT, the goal is to find the minimum number of queries that would allow to identify the defective set, which corresponds to the goal of the attacker in identifying the set $S$, which as few queries to the detector as possible.
@@
1.3: Theoretical Query Complexity. Theoretical analysis of NATGT provides an upper bound on query complexity $T$ to recover $S$:
@@
as derived from [31]. For example, with $d=3\cdot 256^{2}$, $l=32$, and $k=2$, this yields $T=\mathcal{O}(10^{5})$ queries. Such query volume is already impractical if the verifier enforces strict budgets for authenticated accounts to the system, and API restrictions like throttling queries per user.
Step 2: Recover the vector $s$. Even if the adversary succeeds in recovering the unordered set $S$, the attack remains ineffective without knowledge of the ordered fingerprint vector $s$. This is because the reconstructor $R$ is trained to predict fingerprint values at positions specified by the exact ordering of $s$; any mismatch prevents training a surrogate $\hat{R}$ that can faithfully mimic $R$ on out-of-distribution inputs. The overall query complexity is therefore the cost of recovering $S$ in Step 1 plus the cost of recovering the ordering in Step 2.
Recovering $s$ from $S$ requires brute-force search over all permutations of $S$, which has factorial complexity $O(l!)$. For example, for a fingerprint length of $l=32$, this corresponds to approximately $O(10^{35})$ permutations. Thus, for sufficiently large fingerprints, the complexity of Step 2 quickly becomes prohibitive and typically dominates the attack cost, far exceeding that of Step 1. As shown in our experiments with StyleGAN2 and SD on $\times256$ RGB images, the optimal fingerprint length is at least $l=16$, already rendering the attack impractical for any realistic adversary. Attacks against models for small images, such as (e.g., $\times8$), where the complexity of both steps may be feasible; however, such resolutions are unrealistic for modern generative models, which continue to strive toward higher image dimensions.
Ensemble Robustness. Robustness against fingerprint recovery can be further improved via an ensemble of detectors. If the verifier accepts an image when any detector in the ensemble accepts it, the adversary must avoid triggering all detectors simultaneously to succeed. This ensemble strategy increases robustness by reducing FNR and increasing the minimum number of overlapping pixels needed for rejection across detectors. Note that this stands in contrast to max-MSE ensembles aimed at minimizing FPR.
VI Discussion and Conclusion
As generative models are increasingly deployed in high-stakes domains such as healthcare and the legal system, ensuring accountability and maintaining the integrity of certified models after deployment becomes critical. Regulatory frameworks, such as the EU AI Act, underscore the importance of independent audits of machine learning systems. Yet, existing approaches often fail to address the risk of a provider substituting a certified model with a cheaper or lower-quality alternative post-certification. This work has introduced AuthPrint, a passive black-box fingerprinting framework designed to verify whether an image originates from a certified generative model, even in settings where the provider may act adversarially.
VI-A Merits of AuthPrint
AuthPrint advances the state of the art along several dimensions. First, it directly addresses the open-world binary verification problem: determining whether an image was generated by a specific model, rather than classifying among a fixed set of known models, as in prior fingerprinting approaches. This distinction makes it suitable for real-world audit scenarios in which the integrity of a single certified model must be preserved. Second, it operates in a fully black-box setting, requiring no access to model internals, specialized hardware, or parameter modifications (as typically required in watermarking), thereby ensuring broad deployability. Third, existing fingerprinting techniques assume honest model providers and focus on preventing fingerprint removal by external attackers. In contrast, AuthPrint addresses a threat model where the model provider itself may attempt to deceive the fingerprinter—a threat model that had not yet been explored in prior work. To defend against a malicious model provider, AuthPrint employs output-dependent fingerprints tied to secret pixel locations known only to the verifier. By leveraging principles from out-of-distribution detection, this design captures subtle distributional shifts caused by model substitution, while the secrecy of fingerprint indices provides robustness against adaptive adversarial strategies, including evasion and reverse-engineering.
We evaluate AuthPrint on two widely used model families—StyleGAN2 and SD—under diverse substitution strategies, including optimization techniques like quantization, pruning, and post-processing. Experiments show that AuthPrint achieves perfect detection performance on GAN models, with FPR@95%TPR reduced to zero. Diffusion models present additional challenges due to their iterative denoising process; however, we demonstrate that practical strategies, such as scaling training data and reconstructor capacity, and employing ensemble detectors, yield high detection performance on SD models, comparable to StyleGAN2.
From a security perspective, AuthPrint exhibits strong resilience against adaptive adversarial attacks. Gradient-based evasion techniques (e.g., PGD) are ineffective because adversaries lack access to the reconstructor and fingerprint indices. For fingerprint recovery, we reduce the task to group testing and show that recovering the complete set of fingerprinted pixels requires an infeasible number of queries; moreover, reconstructing their precise ordering is even more difficult due to the high query complexity of the attack strategy. These results highlight the robustness of AuthPrint when its internal details remain hidden from the provider.
A caveat lies in the secrecy assumptions underlying $R$. While secrecy of the reconstructor is central to AuthPrint, the practical feasibility of upholding this assumption deserves scrutiny. Our prototype uses a 32M-parameter reconstructor for StyleGAN2 and a 674M-parameter reconstructor for SD, but larger reconstructors may be required to verify high-capacity generators such as the latest SD under diverse prompts. Larger models raise operational risks (e.g., expanded hardware attack surfaces or potential leakage of parameters through infrastructure vulnerabilities) as well as scalability challenges, since the auditor must maintain a distinct $R$ for each certified model. These considerations do not undermine the conceptual validity of AuthPrint, but they highlight deployment trade-offs and the need for trusted, well-resourced third-party verifiers capable of safeguarding $R$ throughout the model’s lifetime.
VI-B Limitations of AuthPrint
Nonetheless, certain limitations remain. Detection accuracy may degrade with highly diverse prompts in conditional models such as SD, and quantization is more difficult to detect than other compression techniques such as pruning. These challenges may be mitigated through improved computational strategies during reconstructor training or through pipeline optimizations, pointing to an important avenue for future research. Extensions of AuthPrint could also target other modalities—including audio, video, and text—or explore fingerprinting in alternative domains such as frequency or hybrid spatial–frequency representations. Finally, integrating cryptographic techniques, such as zero-knowledge proofs or trusted execution environments, could enable powerful hybrid verification frameworks.
VI-C Conclusion
In conclusion, AuthPrint provides the first black-box ML-based method for verifying whether generated images originate from a certified model under a threat model that treats the provider as potentially adversarial. It addresses the open-world verification setting, requires no access to model internals, and demonstrates robustness against a range of adaptive attack strategies. Our evaluation across GAN and diffusion models highlights both the effectiveness of the approach and the remaining challenges, such as diverse prompts and quantization. Overall, AuthPrint offers a promising foundation for auditing generative models and motivates further work on extending verification across modalities and strengthening robustness in real-world deployment scenarios.
VII Usage of LLMs
LLMs were used for editorial purposes in this manuscript, and all outputs were inspected by the authors to ensure accuracy and originality.